# Adversarial Score Identity Distillation: Rapidly Surpassing the Teacher in One Step

### v2

@@ 5 Experimental Results
We evaluate SiDA through comprehensive experiments designed to assess its effectiveness compared to the teacher diffusion model and other existing baseline methods for both unconditional and label-conditional image generation. We present results that measure both quantitative and qualitative performance across various datasets, aiming to demonstrate SiDA’s capability to generate high-quality samples through efficient score distillation. Detailed descriptions of the experimental setup, datasets, and evaluation metrics are provided below.
Datasets and Teacher Models. To comprehensively evaluate the effectiveness of SiDA, we first utilize four representative datasets of varying scales and resolutions and the EDM diffusion models pretrained on them, as discussed in Karras et al. (2022). These datasets include CIFAR-10 ($32\times 32$; both conditional and unconditional) (Krizhevsky et al., 2009), ImageNet ($64\times 64$) (Deng et al., 2009), FFHQ ($64\times 64$) (Karras et al., 2019), and AFHQ-v2 ($64\times 64$) (Choi et al., 2020).
Additionally, we distill EDM2 diffusion models of varying sizes that were pretrained on ImageNet ($512\times 512$), following details provided by Karras et al. (2024). We compare these distilled models to their corresponding teacher models and to a concurrent work of Lu & Song (2024) on EDM2 distillation. The models range from approximately 100 million to 1 billion parameters. By utilizing these diverse datasets and model scales, we thoroughly assess SiDA’s performance across different content types and complexities, ensuring a robust evaluation of its generative capabilities.
Evaluation protocol. We assess the quality of image generation using the Fréchet Inception Distance (FID) and Inception Score (IS) (Salimans et al., 2016). Following Karras et al. (2024), we also evaluate models trained on ImageNet ($512\times 512$) using $\text{FD}_{\text{DINOv2}}$, the Fréchet distance computed in the feature space of DINOv2 (Oquab et al., 2023). Consistent with the methodology of Karras et al. (2019; 2022; 2024), these scores are calculated using 50,000 generated samples, with the training dataset used by the EDM or EDM2 teacher model [^1] https://github.com/NVlabs/edm serving as the reference. Additionally, we evaluate SiDA on ImageNet $64\times 64$ using the Precision and Recall metrics (Kynkäänniemi et al., 2019), where both metrics are calculated using a predefined reference batch [^2] https://openaipublic.blob.core.windows.net/diffusion/jul-2021/ref_batches/imagenet/64/VIRTUAL_imagenet64_labeled.npz [^3] https://github.com/openai/guided-diffusion/tree/main/evaluations (Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Song et al., 2023; Song & Dhariwal, 2023). In line with SiD, we periodically evaluate the FID during distillation and preserve the generators with the lowest FID. To ensure accuracy, we conduct re-evaluations across 10 independent runs to obtain reliable metrics. We recommend this more statistically rigorous approach over reporting only the best metrics observed across independent runs or within the distillation process.
Implementation details. We implement SiDA using the SiD codebase (Zhou et al., 2024b), incorporating necessary functions adapted from either EDM (Karras et al., 2022) or EDM2 (Karras et al., 2024). The score estimation network, $f_{\psi}$, is initialized using the same architecture and parameters as the pre-trained teacher score network, $f_{\phi}$, from either EDM or EDM2. For the initialization of the student generator, $G_{\theta}$, we explore two settings: one where $G_{\theta}$ is initialized with the same parameters as the pre-trained EDM or EDM2 teacher score network, and another where it is initialized using the distilled generator from SiD, referred to as SiD-SiDA (SiD2A).
Our training process unfolds in three stages: 1) For the first 100k images, we exclusively train $f_{\psi}$ to stabilize the score estimation. 2) For the subsequent 100k images, both $f_{\psi}$ and $G_{\theta}$ are trained using the SiD generator loss (6), integrating the distilled insights. 3) For all remaining images, we update $f_{\psi}$ and $G_{\theta}$ with the SiDA generator loss (8), which incorporates adversarial adjustments to refine the generation process. Throughout the training, the score estimation network $f_{\psi}$ is consistently trained with the SiDA fake score loss (9), ensuring that the adversarial components are properly integrated. Additional implementation details are provided in Appendix B.
We observe that using only the adversarial loss, without the SiD component in the combined loss, leads to rapid divergence, even when the single-step generator is initialized from SiD.
Figure 1: Ablation Study of $\alpha$ on distilling an EDM model pretrained on CIFAR-10 (unconditional): Each plot illustrates the relation between the performance, measured by FID (left) and Inception Score (right) vs. the number of training iterations during the distillation, across varying values of $\alpha$. The batch sizer is 256. The study underscores the impact of $\alpha$ on both training efficiency and generative fidelity, leading us to select $\alpha=1.0$ for the subsequent experiments.
Figure 2: Analogous plot for SiD2A, a SiD initialized SiDA model. We observe a trajectory of improvements in SiD2A over the SiD-distilled generator following an initial warmup period, and more robust performance in terms of the choice of $\alpha$.
@@
We conducted an ablation study to evaluate the impact of the gradient bias correction weight factor $\alpha$, a pivotal hyperparameter for SiD, on the performance of SiDA. Specifically, we tested a spectrum of $\alpha$ values $[-0.25,0.0,0.5,0.75,1.0,1.2,1.5]$ during the distillation of the EDM model pretrained on CIFAR-10 in an unconditional context. Figure 2 demonstrates how the model’s performance varies with different $\alpha$ values over the first 20k iterations at a batch size of 256, assessed using FID and IS. We determined that an $\alpha$ value of 1.0 yielded the best performance, prompting us to adopt this setting for subsequent experiments of SiDA over various datasets. In contrast, lower $\alpha$ values such as $-0.25$ or 0 resulted in suboptimal performance, while high values like 1.2 or 1.5 led to training instability and divergence. Interestingly, the performance disparity for $\alpha$ values between 0.5 and 1.0 was significantly narrower than those observed in SiD, as depicted in Figure 8 of Zhou et al. (2024b). These findings imply that although the adversarial loss may introduce some instability, it could also compensate for the inadequate correction of gradient biases associated with smaller $\alpha$ values.
@@
In an alternative setting, referred to as SiD2A (SiD-initialized SiDA), where the generator $G_{\theta}$ is initialized using the best available pre-distilled SiD generator, we observe robust and stabilized training across all choices of $\alpha$. This suggests that initializing with the pre-distilled SiD generator provides a more advantageous starting point to achieve enhanced training stability and generative quality. Following these observations, and given that $\alpha=1.0$ and $\alpha=1.2$ were also utilized in SiD experiments, we have selected these same values for subsequent experiments with SiD2A.
@@
Family Model NFE FID ($\downarrow$) IS ($\uparrow$) Teacher VP-EDM (Karras et al., 2022) 35 1.97 9.68 Diffusion DDPM (Ho et al., 2020) 1000 3.17 9.46 $\pm$ 0.11 DDIM (Song et al., 2020) 100 4.16 DPM-Solver-3 (Lu et al., 2022) 48 2.65 VDM (Kingma et al., 2021) 1000 4.00 iDDPM (Nichol & Dhariwal, 2021) 4000 2.90 HSIVI-SM (Yu et al., 2023) 15 4.17 TDPM+ (Zheng et al., 2022) 100 2.83 9.34 VP-EDM+LEGO-PR (Zheng et al., 2024) 35 1.88 9.84 One Step StyleGAN2+ADA+Tune (Karras et al., 2020) 1 2.92 $\pm$ 0.05 9.83 $\pm$ 0.04 Diffusion ProjectedGAN (Wang et al., 2023b) 1 2.54 iCT-deep (Song & Dhariwal, 2023) 1 2.51 9.76 Diff-Instruct (Luo et al., 2023b) 1 4.53 9.89 StyleGAN2+ADA+Tune+DI (Luo et al., 2023b) 1 2.71 9.86 $\pm$ 0.04 DMD (Yin et al., 2024b) 1 3.77 CTM (Kim et al., 2023) 1 1.98 GDD-I (Zheng & Yang, 2024) 1 1.54 10.10 SiD, $\alpha=1.0$ 1 2.028 $\pm$ 0.020 10.017 $\pm$ 0.047 SiD, $\alpha=1.2$ 1 1.923 $\pm$ 0.017 9.980 $\pm$ 0.042 SiDA, $\alpha=1.0$ 1 1.516 $\pm$ 0.010 10.323 $\pm$ 0.048 SiD2A, $\alpha=1.0$ 1 1.499 $\pm$ 0.012 10.188 $\pm$ 0.035 SiD2A, $\alpha=1.2$ 1 1.519 $\pm$ 0.009 10.252 $\pm$ 0.027
@@
Table 3: Analogous to Table 2 for ImageNet 64x64 with label conditioning. The Precision and Recall metrics are also included.
Family Model NFE FID ($\downarrow$) Prec. ($\uparrow$) Rec. ($\uparrow$) Teacher VP-EDM (Karras et al., 2022) 511 1.36 79 2.64 0.71 0.67 Diffusion RIN (Jabri et al., 2022) 1000 1.23 DDPM (Ho et al., 2020) 250 11.00 0.67 0.58 ADM (Dhariwal & Nichol, 2021) 250 2.07 0.74 0.63 DPM-Solver-3 (Lu et al., 2022) 50 17.52 HSIVI-SM (Yu et al., 2023) 15 15.49 U-ViT (Bao et al., 2022) 50 4.26 DiT-L/2 (Peebles & Xie, 2023) 250 2.91 LEGO (Zheng et al., 2024) 250 2.16 Consistency iCT (Song & Dhariwal, 2023) 1 4.02 0.70 0.63 iCT-deep (Song & Dhariwal, 2023) 1 3.25 0.72 0.63 Distillation PD (Salimans & Ho, 2022) 2 8.95 0.63 0.65 G-distill (Meng et al., 2023) ($w$=0.3) 8 2.05 BOOT (Gu et al., 2023) 1 16.3 0.68 0.36 PID (Tee et al., 2024) 1 9.49 DFNO (Zheng et al., 2023) 1 7.83 0.61 CD-LPIPS (Song et al., 2023) 2 4.70 0.69 0.64 Diff-Instruct (Luo et al., 2023b) 1 5.57 TRACT (Berthelot et al., 2023) 2 4.97 DMD (Yin et al., 2024b) 1 2.62 CTM (Kim et al., 2023) 1 1.92 0.57 CTM (Kim et al., 2023) 2 1.73 0.57 GDD-I (Zheng & Yang, 2024) 1 1.16 0.75 0.60 EMD-16 (Xie et al., 2024) 1 2.2 0.59 DMD2  (Yin et al., 2024a) 1 1.51 DMD2+longer training  (Yin et al., 2024a) 1 1.28 SiD , $\alpha=1.0$ 1 2.022 $\pm$ 0.031 0.73 0.63 SiD , $\alpha=1.2$ 1 1.524 $\pm$ 0.009 0.74 0.63 SiDA, $\alpha=1.0$ (ours) 1 1.353 $\pm$ 0.025 0.74 0.63 SiD2A, $\alpha=1.0$ (ours) 1 1.114 $\pm$ 0.019 0.75 0.62 SiD2A, $\alpha=1.2$ (ours) 1 1.110 $\pm$ 0.018 0.75 0.62
@@ 5.2 Benchmarking EDM Distillation
We begin by assessing the performance of SiDA and SiD2A, focusing on both generation quality and convergence speed across the datasets previously mentioned. Our comprehensive evaluation compares SiDA and SiD2A against top-performing deep generative models, including GANs, traditional diffusion models, and their distilled counterparts, in terms of generation quality. Qualitatively, random images generated by SiD2A in a single step are displayed from Figures 6 to 10 in the Appendix, illustrating the high quality achieved.
@@
We further validate the effectiveness of SiDA and SiD2A on the ImageNet 64x64 dataset. Here, SiDA demonstrated superior performance over SiD, achieving an FID of 1.353. Notably, by leveraging the advanced capabilities of the SiD one-step generator, SiD2A attained an unprecedented low FID of 1.110 on ImageNet 64x64 with $\alpha=1.2$, surpassing the teacher model’s performance by a large margin. On the FFHQ 64x64 and AFHQ-v2 64x64 datasets, known for human and animal faces, SiDA and SiD2A again prove superior. The datasets’ less diverse patterns do not diminish the models’ performance, with SiDA and SiD2A maintaining faster convergence and surpassing the teacher model rapidly. The clear improvements across diverse datasets underscore SiDA and SiD2A’s effectiveness.
It’s worth noting that SiDA’s close competitor for ImageNet ($64\times 64$), GDD-I (Zheng & Yang, 2024), uses a discriminator inspired by Projected GAN (Sauer et al., 2021), integrating features from VGG16-BN and EfficientNet-lite0. Concerns have been raised, as discussed in Kynkäänniemi et al. (2023) and by Song & Dhariwal (2023), about the potential for encoders pretrained on ImageNet to inadvertently enhance FID scores through feature leakage. This potential issue does not affect SiD, SiDA, or SiD2A, ensuring our results reflect true model performance without such confounds.
SiDA also demonstrates accelerated convergence, as evidenced in Figures 3 and 5. The logarithmic scale plots show SiDA’s FID decreasing more rapidly than SiD’s, an efficiency that also holds when scaled up to ImageNet 64x64. Overall, SiDA consistently outperforms SiD in terms of convergence speed by a substantial margin, marking a significant acceleration, particularly noteworthy since SiD already converges at an exponential rate. This rapid improvement in both speed and performance underscores SiDA’s effectiveness and efficiency in generating high-quality images across different datasets. These results not only highlight SiD2A’s enhanced image quality and reduced iteration needs but also the overall efficiency of SiDA and SiD2A across various datasets, confirming the advantages of these advanced distillation techniques.
@@ 5.3 Benchmarking EDM2 Distillation
We validate the effectiveness of SiD, SiDA, and SiD2A in distilling EDM2 models of six different sizes pretrained on ImageNet 512x512. Our findings are summarized in Tables 4 and 7 and illustrated in Figure 4. Notably, SiD, our baseline single-step generator introduced six months ago on arXiv, performs comparably to the EDM2 teacher, which requires 63 NFEs, as well as to the simple, stable, and scalable consistency distillation (sCD) method of Lu & Song (2024) that uses 2 NFEs—a concurrent work to SiDA. SiD also convincingly outperforms sCD with 1 NFE, despite not yet employing CFG during distillation. Consistent with findings from EDM distillation, both SiDA and SiD2A demonstrate superior performance over SiD in distilling EDM2, achieving FIDs of 1.756 and 1.669, respectively, using the EDM2-S model with only 280M parameters. These results already surpass the EDM2-XXL model and sCD at 1.5B parameters, whose FIDs exceed 1.8.
Notably, scaling SiDA to EDM2-XL achieves a record-low FID of 1.449. While results for SiD2A on EDM2-XL and for the SiD family on EDM2-XXL are forthcoming, Figure 4 indicates the potential for further FID reductions as these results become available.
Mirroring findings from EDM distillation, SiDA also exhibits accelerated convergence when distilling EDM2, as shown in Figure 4. Logarithmic-scale plots reveal a more rapid decrease in FID for SiDA compared to SiD, an efficiency that scales effectively to ImageNet 512x512 and across generators of various sizes. Overall, SiDA consistently surpasses prior and concurrent methods in terms of both final metrics and convergence speed. This rapid improvement highlights SiDA’s effectiveness and efficiency in generating high-resolution images across various model sizes.
Table 4: FID scores and number of parameters for different methods on ImageNet (512x512).
@@
SiD N 1 3.353 $\pm$ 0.041 2.707 $\pm$ 0.054 2.060 $\pm$ 0.038 1.907 $\pm$ 0.016 2.005 $\pm$ 0.024 -
SiDA N 1 2.228 $\pm$ 0.037 1.756 $\pm$ 0.023 1.546 $\pm$ 0.023 1.501 $\pm$ 0.024 1.449 $\pm$ 0.033 -
SiD2A N 1 2.156 $\pm$ 0.028 1.669 $\pm$ 0.019 1.488 $\pm$ 0.038 1.465 $\pm$ 0.028 - -
(a) EDM-XS
(b) EDM-S
(c) EDM-M
(d) EDM-L
(e) EDM-XL
(f) EDM-XXL
@@
Figure 4: Evolution of FIDs for the SiD and SiDA generators during the distillation of the EDM2 teacher models of six different sizes pretrained on ImageNet (512x512), with a batch size of 2048 and $\alpha=1.0$. The performance of EDM2 without classifier-free guidance (CFG) and EDM2 with CFG, using 64 NFEs, along with the simple, stable, and scalable consistency distillation (sCD) method of Lu & Song (2024) with 1 and 2 NFEs, is depicted with horizontal lines in purple, red, green, and cyan respectively.

### v4

@@ 5 Experimental Results
We evaluate SiDA through comprehensive experiments designed to assess its effectiveness compared to the teacher diffusion model and other existing baseline methods for both unconditional and label-conditional image generation. We present results that measure both quantitative and qualitative performance across various datasets, aiming to demonstrate SiDA’s capability to generate high-quality samples through efficient score distillation.
We focus on distilling EDM (Karras et al., 2022) diffusion models pretrained on four distinct datasets, as well as six different-sized EDM2 (Karras et al., 2024), all pretrained on ImageNet 512 $\times$ 512. Notably, while the SiD technique (Zhou et al., 2024b) has been evaluated using EDM, it has not yet been tested on EDM2. This paper makes a supplementary contribution by adapting the SiD codebase for compatibility with the EDM2 framework, which represents an advanced evolution of EDM, and establishing it as a stong baseline for comparison. For detailed descriptions of the experimental setup, datasets, and evaluation metrics, please refer to Appendix A. For details of the algorithm and modifications needed to adapt SiD and SiDA from EDM to EDM2, prlease refer to Appendix B.
As outlined in Tables 6 and 7, SiDA introduces negligible memory overhead compared to SiD, owing to its design that introduces no additional parameters. Although SiDA experiences an approximate 10% reduction in iteration speed and requires real data, it converges significantly faster than SiD and consistently delivers superior performance relative to the teacher model, as demonstrated below.
@@
Figure 2: Ablation Study of $\alpha$ on distilling an EDM model pretrained on CIFAR-10 (unconditional): Each plot illustrates the relation between the performance, measured by FID (left) and Inception Score (right) vs. the number of training iterations during distillation, across varying values of $\alpha$. The batch sizer is 256. The study underscores the impact of $\alpha$ on both training efficiency and generative fidelity, leading us to select $\alpha=1.0$ for subsequent experiments.
Figure 3: Analogous plot for SiD2A, a SiD initialized SiDA model. We observe a trajectory of improvements in SiD2A following an initial warmup period, and more robust performance relative to $\alpha$ choice.
@@
The gradient bias correction weight factor $\alpha$ is an important hyperparameter in SiD, typically set to 1 or 1.2. We conducted an ablation study to assess its impact on the performance of SiDA. Specifically, we tested a spectrum of $\alpha$ values $[-0.25,0.0,0.5,0.75,1.0,1.2,1.5]$ during the distillation of the EDM model pretrained on CIFAR-10 (unconditional). Figure 3 demonstrates how the model’s performance varies with different $\alpha$ values over the first 20k iterations at a batch size of 256, assessed using FID and IS. We determined that an $\alpha$ value of 1.0 yielded the best performance, prompting us to adopt this setting for subsequent experiments of SiDA over various datasets. In contrast, lower $\alpha$ values such as $-0.25$ or 0 resulted in suboptimal performance, while high values like 1.2 or 1.5 led to training instability and divergence. Interestingly, the performance disparity for $\alpha$ values between 0.5 and 1.0 was significantly narrower than those observed in SiD, as depicted in Figure 8 of Zhou et al. (2024b). These findings imply that although the adversarial loss may introduce some instability into the distillation process, it could also compensate for the inadequate correction of gradient biases associated with smaller $\alpha$ values.
@@
In an alternative setting, referred to as SiD2A (SiD-initialized SiDA), where the generator $G_{\theta}$ is initialized using the best available pre-distilled SiD generator, we observe robust and stabilized training across all choices of $\alpha$. This suggests that initializing with the pre-distilled SiD generator provides a more advantageous starting point to achieve enhanced training stability and generative quality. Following these observations, and given that $\alpha=1.0$ and $\alpha=1.2$ were also utilized in SiD experiments, we have selected these same values for subsequent experiments with SiD2A for EDM.
@@
Family Model NFE FID ($\downarrow$) IS ($\uparrow$) Teacher VP-EDM (Karras et al., 2022) 35 1.97 9.68 Diffusion DDPM (Ho et al., 2020) 1000 3.17 9.46 $\pm$ 0.11 DDIM (Song et al., 2020) 100 4.16 DPM-Solver-3 (Lu et al., 2022) 48 2.65 VDM (Kingma et al., 2021) 1000 4.00 iDDPM (Nichol & Dhariwal, 2021) 4000 2.90 HSIVI-SM (Yu et al., 2023b) 15 4.17 TDPM+ (Zheng et al., 2022) 100 2.83 9.34 VP-EDM+LEGO-PR (Zheng et al., 2024b) 35 1.88 9.84 One Step StyleGAN2+ADA+Tune (Karras et al., 2020) 1 2.92 $\pm$ 0.05 9.83 $\pm$ 0.04 Diffusion ProjectedGAN (Wang et al., 2023b) 1 2.54 iCT-deep (Song & Dhariwal, 2023) 1 2.51 9.76 Diff-Instruct (Luo et al., 2023b) 1 4.53 9.89 StyleGAN2+ADA+Tune+DI (Luo et al., 2023b) 1 2.71 9.86 $\pm$ 0.04 DMD (Yin et al., 2024b) 1 3.77 CTM (Kim et al., 2023) 1 1.98 GDD-I (Zheng & Yang, 2024) 1 1.54 10.10 SiD, $\alpha=1.0$ 1 2.028 $\pm$ 0.020 10.017 $\pm$ 0.047 SiD, $\alpha=1.2$ 1 1.923 $\pm$ 0.017 9.980 $\pm$ 0.042 SiDA, $\alpha=1.0$ 1 1.516 $\pm$ 0.010 10.323 $\pm$ 0.048 SiD2A, $\alpha=1.0$ 1 1.499 $\pm$ 0.012 10.188 $\pm$ 0.035 SiD2A, $\alpha=1.2$ 1 1.519 $\pm$ 0.009 10.252 $\pm$ 0.027
@@ 5.2 Benchmarking EDM Distillation
We begin by assessing the performance of SiDA and SiD2A, focusing on both generation quality and convergence speed across different datasets. Our comprehensive evaluation compares SiDA and SiD2A to leading deep generative models, including GANs, diffusion models, their distilled counterparts, and various variations, focusing on generation quality. Qualitatively, random images generated by SiD2A in a single step are displayed from Figures 11 to 15 in the Appendix.
@@
We further validate the effectiveness of SiDA and SiD2A on the ImageNet 64x64 dataset. Here, SiDA demonstrated superior performance over SiD, achieving an FID of 1.353. Notably, by leveraging the advanced capabilities of the SiD one-step generator, SiD2A attained an unprecedented low FID of 1.110 on ImageNet 64x64 with $\alpha=1.2$, surpassing the teacher model’s performance by a large margin. On the FFHQ 64x64 and AFHQ-v2 64x64 datasets, known for human and animal faces, SiDA and SiD2A again prove superior. The datasets’ less diverse patterns do not diminish the models’ performance, with SiDA and SiD2A maintaining
Table 3: Analogous to Table 2 for ImageNet 64x64 with label conditioning, using a pretrained EDM diffusion model as the teacher. The Precision and Recall metrics are also included.
Family Model NFE FID ($\downarrow$) Prec. ($\uparrow$) Rec. ($\uparrow$) Teacher VP-EDM (Karras et al., 2022) 511 1.36 79 2.64 0.71 0.67 Diffusion RIN (Jabri et al., 2022) 1000 1.23 DDPM (Ho et al., 2020) 250 11.00 0.67 0.58 ADM (Dhariwal & Nichol, 2021a) 250 2.07 0.74 0.63 DPM-Solver-3 (Lu et al., 2022) 50 17.52 HSIVI-SM (Yu et al., 2023b) 15 15.49 U-ViT (Bao et al., 2022) 50 4.26 DiT-L/2 (Peebles & Xie, 2023) 250 2.91 LEGO (Zheng et al., 2024b) 250 2.16 Consistency iCT (Song & Dhariwal, 2023) 1 4.02 0.70 0.63 iCT-deep (Song & Dhariwal, 2023) 1 3.25 0.72 0.63 Distillation PD (Salimans & Ho, 2022) 2 8.95 0.63 0.65 G-distill (Meng et al., 2023) ($w$=0.3) 8 2.05 BOOT (Gu et al., 2023) 1 16.3 0.68 0.36 PID (Tee et al., 2024) 1 9.49 DFNO (Zheng et al., 2023) 1 7.83 0.61 CD-LPIPS (Song et al., 2023) 2 4.70 0.69 0.64 Diff-Instruct (Luo et al., 2023b) 1 5.57 TRACT (Berthelot et al., 2023) 2 4.97 DMD (Yin et al., 2024b) 1 2.62 CTM (Kim et al., 2023) 1 1.92 0.57 CTM (Kim et al., 2023) 2 1.73 0.57 GDD-I (Zheng & Yang, 2024) 1 1.16 0.75 0.60 EMD-16 (Xie et al., 2024) 1 2.2 0.59 DMD2  (Yin et al., 2024a) 1 1.51 DMD2+longer training  (Yin et al., 2024a) 1 1.28 SiD , $\alpha=1.0$ 1 2.022 $\pm$ 0.031 0.73 0.63 SiD , $\alpha=1.2$ 1 1.524 $\pm$ 0.009 0.74 0.63 SiDA, $\alpha=1.0$ (ours) 1 1.353 $\pm$ 0.025 0.74 0.63 SiD2A, $\alpha=1.0$ (ours) 1 1.114 $\pm$ 0.019 0.75 0.62 SiD2A, $\alpha=1.2$ (ours) 1 1.110 $\pm$ 0.018 0.75 0.62
faster convergence and surpassing the teacher model rapidly. The clear improvements across diverse datasets underscore their effectiveness.
It’s worth noting that SiDA’s close competitor for ImageNet 64x64, GDD-I (Zheng & Yang, 2024), uses a discriminator inspired by Projected GAN (Sauer et al., 2021), integrating features from VGG16-BN and EfficientNet-lite0. Concerns arose, as discussed in Kynkäänniemi et al. (2023) and Song & Dhariwal (2023), about the potential for encoders pretrained on ImageNet to inadvertently enhance FID scores through feature leakage. This potential issue does not affect SiD, SiDA, or SiD2A, ensuring our results reflect true model performance without such confounds.
SiDA also demonstrates accelerated convergence, as evidenced in Figures 4 and 6. The logarithmic scale plots show SiDA’s FID decreasing more rapidly than SiD’s, an efficiency that also holds when scaled up to ImageNet 64x64. Overall, SiDA consistently outperforms SiD in terms of convergence speed by a substantial margin, marking a significant acceleration, particularly noteworthy since SiD already converges at an exponential rate. This rapid improvement in both speed and performance underscores SiDA’s effectiveness and efficiency in generating high-quality images across different datasets. These results not only highlight SiD2A’s enhanced image quality and reduced iteration needs but also the overall efficiency of SiDA and SiD2A across various datasets, confirming the advantages of these advanced distillation techniques.
@@ 5.3 Benchmarking EDM2 Distillation
Table 4: FID scores and number of parameters for various methods on ImageNet 512x512 are presented. The results for EDM2 are sourced from Karras et al. (2024), while those for sCT and sCD are obtained from Lu & Song (2024).
@@
SiD N 1 3.353 $\pm$ 0.041 2.707 $\pm$ 0.054 2.060 $\pm$ 0.038 1.907 $\pm$ 0.016 1.888 $\pm$ 0.030 1.969 $\pm$ 0.029
SiDA N 1 2.228 $\pm$ 0.037 1.756 $\pm$ 0.023 1.546 $\pm$ 0.023 1.501 $\pm$ 0.024 1.428 $\pm$ 0.018 1.503 $\pm$ 0.020
SiD2A N 1 2.156 $\pm$ 0.028 1.669 $\pm$ 0.019 1.488 $\pm$ 0.038 1.413 $\pm$ 0.022 1.379 $\pm$ 0.017 1.366 $\pm$ 0.015
We evaluate the effectiveness of SiD, SiDA, and SiD2A in distilling EDM2 models of six different sizes, all pretrained on ImageNet 512x512. Our findings and comparisons are detailed in Tables 4, 5, and 10 and depicted in Figure 5. The rapid convergence of the algorithms is visually illustrated in Figures 8 through 10. Random images generated by SiD2A in a single step are displayed in Figures 16 through 34. Notably, SiD, the baseline single-step algorithm on which SiDA is based, already performs comparably to the EDM2 teacher, which requires 63 NFEs, and to the stable and scalable consistency distillation (sCD) method of Lu & Song (2024) that uses 2 NFEs—a concurrent development to SiDA. Moreover, SiD convincingly outperforms sCD with 1 NFE, even without using CFG during distillation. As indicated in Table 5, SiD, operating with a single step and without CFG in a data-free setting, ranks among the top performers of all generative models reported on ImageNet 512x512. This state-of-the-art performance provides a robust foundation for both SiDA and SiD2A, which will incorporate real data in their processes.
In line with these distillation results for the EDM family, both SiDA and SiD2A show superior performance over SiD in distilling EDM2, achieving FIDs of 1.756 and 1.669, respectively, with the EDM2-S model, which has only 280M parameters. These scores already outperform the EDM2-XXL model and sCD at 1.5B parameters, whose FIDs exceed 1.8. Remarkably, when scaled to EDM2-XL and EDM2-XXL, SiD2A achieves record-low FIDs of 1.379 and 1.366, respectively, significantly surpassing previous methods, whether single or multi-step.
Table 5: Performance comparison of generative models trained on ImageNet at 512×512 resolution (Left: without Classifier-Free Guidance (CFG); Right: with CFG). The results for EDM2 of six different sizes and their distillations using sCD, SiD, and SiDA algorithms are provided in Table 4.
Method (CFG=N) NFE (↓) #Params FID (↓)
ADM-G (Dhariwal & Nichol, 2021b) 250 559M 23.24
U-DiT-B (Tian et al., 2024b) 250 204M 15.39
DiT-XL/2 (Peebles & Xie, 2022) 250 675M 12.03
MaskDiT (Zheng et al., 2024a) 79 736M 10.79
ADM-U (Dhariwal & Nichol, 2021b) 250 $>$559M 9.96
LEGO-XL-PR (Zheng et al., 2024b) 250 681M 9.01
BigGAN-deep (Brock et al., 2019) 1 160M 8.43
DiMR-XL/3R (Liu et al., 2024) 250 525M 7.93
MaskGIT (Chang et al., 2022) 12 227M 7.32
MAGVIT-v2 (Yu et al., 2023a) 12 307M 4.61
RIN (Jabri et al., 2022) 1000 320M 3.95
MAGVIT-v2 (Yu et al., 2023a) 64 307M 3.07
VDM++ (Kingma & Gao, 2023) 512 2B 2.99
MAR (Li et al., 2024) 100 481M 2.74
StyleGAN-XL (Sauer et al., 2022) 1×2 168M 2.41
SiD2A-EDM2-XS (ours) 1 125M 2.156
SiD2A-EDM2-S (ours) 1 280M 1.669
SiD2A-EDM2-M (ours) 1 498M 1.488
SiD2A-EDM2-L (ours) 1  [^7] M 1.413
SiD2A-EDM2-XL (ours) 1 1.1B 1.379
SiD2A-EDM2-XXL (ours) 1 1.5B 1.366
Method (CFG=Y) NFE (↓) #Params FID (↓)
ADM-G (Dhariwal & Nichol, 2021b) 250×2 $>$559M 7.72
U-ViT-L/4 (Bao et al., 2023) 250×2 287M 4.67
U-ViT-H/4 (Bao et al., 2023) 250×2 501M 4.05
ADM-U (Dhariwal & Nichol, 2021b) 250×2 $>$559M 3.85
LEGO-XL-PR (Zheng et al., 2024b) 250×2 681M 3.99
DiM-H (Teng et al., 2024) 250×2 860M 3.78
LEGO-XL-PG (Zheng et al., 2024b) 250×2 681M 3.74
DyDiT-XL (Zhao et al., 2024) 100×2 675M 3.61
DiffuSSM-XL-G (Yan et al., 2024) 250×2 660M 3.41
DiT-XL/2 (Peebles & Xie, 2022) 250×2 675M 3.04
DRWKV-H/2 (Fei et al., 2024b) 250×2 779M 2.95
DiMR-XL/3R (Liu et al., 2024) 250×2 525M 2.89
DiS-H/2 (Fei et al., 2024a) 250×2 900M 2.88
DiffiT (Hatamizadeh et al., 2025) 250×2 561M 2.67
VDM++ (Kingma & Gao, 2023) 512×2 2B 2.65
VAR-d36-s (Tian et al., 2024a) 10×2 2.3B 2.63
SiT-XL (Ma et al., 2024) 250×2 675M 2.62
Large-DiT-3B-G [^1] https://github.com/Alpha-VLLM/LLaMA2-Accessory/blob/main/Large-DiT-ImageNet/assets/table.png (Zhang et al., 2023) 250×2 3B 2.52
MaskDiT-G (Zheng et al., 2024a) 79×2 736M 2.50
MAGVIT-v2 (Yu et al., 2023a) 64×2 307M 1.91
MAR [^2] MAR also requires 256 auto-regressive steps before performing reverse diffusion (Li et al., 2024) 100×2 481M 1.73
Mirroring findings from EDM distillation, SiDA also exhibits accelerated convergence when distilling EDM2, as shown in Figure 5. Logarithmic-scale plots show that SiDA achieves an order of magnitude faster decrease in FID compared to SiD, with efficiency that scales effectively to ImageNet 512x512 and across generators of varying sizes. For instance, while SiD required training the generator with approximately 100M generator-synthesized images to match or surpass the teacher’s performance, SiDA needed only 10M such images to exceed the teacher’s performance. Overall, while SiD proves to be a strong baseline, rivaling the performance of the teacher and methods published subsequently, SiDA consistently surpasses both the teacher and SiD by large margins in terms of final metrics and convergence speed. This rapid improvement underscores SiDA’s effectiveness and efficiency in generating high-resolution images in a single step across various model sizes.
(a)      (a) EDM2-XS
(b)      (b) EDM2-S
(c)      (c) EDM2-M
(d)      (d) EDM2-L
(e)      (e) EDM2-XL
(f)      (f) EDM2-XXL
@@
Figure 5: Evolution of FIDs for the SiD, SiDA, and SiD2A generators during the distillation of the EDM2 teacher models of six different sizes pretrained on ImageNet 512x512, with a batch size of 2048 and $\alpha=1.0$. The performance of EDM2 without classifier-free guidance (CFG) and EDM2 with CFG, using 63 NFEs, along with the simple, stable, and scalable consistency distillation (sCD) method of Lu & Song (2024) with 1 and 2 NFEs, is depicted with horizontal lines in purple, red, green, and cyan respectively.